3) Write concisely, using sigma notation:

(a) 2 + 5 + 8 + 11 + 14 + 17

(b) 1 + 5 + 9 + 13 + · · · (to n terms)

(c) (1)(2) + (2)(3) + (3)(4) + (4)(5) + · · · + (n − 1)n

(d) (1)(2)(4) + (2)(3)(6) + (3)(4)(8) + (4)(5)(10) + · · · (to n terms)

(e) 3 + 32 + 33 + 34 + 35 + · · · + 310



1
Expert's answer
2021-05-31T16:04:39-0400

"(a) \newline
2 + 5 + 8 + 11 + 14 + 17+...\newline
=\sum_{n=1}^\infty(3n-1)
\newline
(b)\newline
 1 + 5 + 9 + 13 + · · · (to \space n \space
 terms)\newline
=\sum_{n=1}^n (4n-3)
\newline
(c)\newline
 (1)(2) + (2)(3) + (3)(4) + (4)(5) + · · · + (n − 1)n\newline
=\sum_{n=1}^n n(n+1)
\newline
(d)\newline
 (1)(2)(4) + (2)(3)(6) + (3)(4)(8) + (4)(5)(10) + · · · (to \space
 n \space
 terms)\newline
=\sum_{n=1}^n n(n+1)(2n+2)
\newline
(e)\newline
 3 + 3^2 + 3^3 + 3^4 + 3^5 + · · · + 3^{10}\newline
=\sum_{n=1}^{10} 3^n"